# Dielectric Tensor Prediction for Inorganic Materials Using Latent Information from Preferred Potential

### v1

@@ 2.4.3 Virtual screening & DFPT validation
Figure 6: (a-b) The joint distribution of band gaps $E_{g}$ and electronic dielectric constants $\epsilon^{\infty}$ (a) or ionic dielectric constants $\epsilon^{0}$ (b) for 6,648 training structures and 14,375 stable candidate structures. $\epsilon^{\infty}$ and $\epsilon^{0}$ of training structures and $E_{g}$ of all structures are obtained from computational results in MP, while $\epsilon^{\infty}$ and $\epsilon^{0}$ of candidate structures are predicted by DTNN-$\epsilon^{\infty}/\epsilon^{0}$. The 5-level densities of data points are drawn in contours using the kernel density estimation for structures both in the candidate set and original set. (c) The top 20 most frequently occurring elements in the top 100 high-quality candidate materials screened from 14,375 candidate structures. (d) DFT-calculated $E_{g}-\epsilon$ data points colored by their figure of merit defined by $E_{g}\cdot\epsilon$ for 4,146 stable structures in training data and 12 new stable structures screened by DTNN. (e) Structure visualizaiton of Ba2SmTaO6 as the top-tier candidate validated by DFPT calculation.
@@ 3 Discussion
Transfer learning is the most popular approach to tackle the data bottleneck which is common in materials science. Previous studies have utilized transfer learning in GNN models, both for the same properties and across different properties, to demonstrate enhanced accuracy in materials science[44, 45, 15]. For example, Chen et al. developed the AtomSets framework[46], which employs a straightforward model architecture to process structure-embedded scalar information read from a pretrained MEGNet model[25], enabling high accuracy predictions with limited datasets. However, the feasibility of transfer learning for properties across different tensorial orders has been less explored. Our study indicates that more than scalar features, it is also possible to transfer higher-rank equivariant compositional and structural information from a pretrained EGNN energy model to a different tensorial property prediction with a small amount of available data, while maintaining the equivariance. These equivariant representations may hold shared structural or even electronic-level information, since the TeaNet/PFP model aims to emulate iterative electronic relaxation. This is evident through the accurate predictions our model achieved for both electronic and ionic dielectric constants, which are produced by different microscopic mechanisms. Furthermore, the pretrained TeaNet/PFP serves as an encoder that integrates electron density information, assisting in comprehending the electronic-level relationship between the $E_{g}$ and $\epsilon^{\infty}$. Apart from dielectric constants, several works attempted to predict tensorial properties of materials such as the polarizability[47], quadrupole[48], by using equivariant architectures. The accuracy of such tasks could potentially be improved by fixing parameters in the shallower layers of an EGNN model pretrained on commonly computed properties like energy.
@@
The PFP model is pretrained on over 22 million structures covering the diverse chemical space. It is efficient and expressive to take a a pretrained PFP as the parent model to generate locally-interacted features as the input to a much simpler child model. In the current readout architecture, it consists of only 0.6 million trainable parameters for the downstream tasks of dielectric constant prediction. The model performing best on the validation accuracy appears at the average epoch of 263 out of 5 training runs for total dielectric constants. By stopping the training after 200 epochs without improvement in validation accuracy, the model training can be efficiently completed within approximately 2 hours using about 6GB memory on an Nvidia Tesla-V100 GPU. We proved that such efficient training on encodings from the parent model can bring a promising property predictor even on a small data region as demonstrated in Table 2. The incorporation of second-order equivariant features in our model may also contribute to high data efficiency as the symmetry of the dielectric tensors is automatically guaranteed due to the model architecture. Based on only 6.6k sample structures with labeled tensor containing up to 6 independent elements from MP, our model achieved a mean tensor prediction error of $0.368$ for $\boldsymbol{\epsilon}^{\infty}$, $1.677$ for $\boldsymbol{\epsilon}^{0}$ and $1.911$ for $\boldsymbol{\epsilon}$.
@@
After validating the screening results with DFPT computations, our model successfully identified Ba2SmTaO6 out of massive data with over 14k structures directly based on its structural graph. Thanks to widely-supported elements in the pretrained PFP, our model can also give relatively good prediction on rare earth elements, e.g., the Sm element in Ba2SmTaO6. The rare earth elements have diverse industrial applications such as glass, lights, magnets, batteries, and catalytic converters and so on[49], due to their unique physical and chemical properties. Our model has the potential to make predictions about properties concerning rare elements, which typically suffer from limited data availability.
@@
In this study, we shows the applicability of our model to a problem in microelectronic device manufacturing, where only the simplified scalar dielectric constant is taken to evaluate the charge storage ability. Accurate predictions of dielectric constants as tensors can contribute to understanding material behavior in the presence of electric fields and their interactions with other materials. This knowledge can facilitate the development of new materials with tailored properties for specific applications, such as energy storage, signal transmission, and electronic component design.
@@ 4 Conclusion
An equivariant model is introduced to predict the second-order tensorial property, dielectric constants, for inorganic materials across 72 supported elements. By leveraging transfer learning, the TeaNet/PFP was treated as the frozen base model to encode material graphs with integrated elemental and structural information. A light-weight trainable equivariant readout module is connected to integrate information in different orders for the final outputs. It was discovered that embedded features from deeper GCN layers in TeaNet/PFP contribute to better accuracy in predicting total dielectric constants. After efficiently training the equivariant readout module, our DTNN model achieved higher accuracy in predicting the electronic, ionic, total dielectric tensors compared with state-of-the-art models, PaiNN, M3GNet and MatTen, across various crystal system types. To test the model’s capability in discovering novel dielectric materials for microelectronic manufacturing with large band gaps and high dielectric constants, we observed that our model can comprehend the electronic-level inverse correlation between band gaps and electronic dielectric constants. Finally, the virtual screening conducted by the model identified one dielectric material that contains a combination of high-dielectric constant and large band gap, Ba2SmTaO6 ($E_{g}=3.36\mathrm{eV},\epsilon=93.81$) from a vast pool of over 14k candidates with only 12 DFPT calculations.
@@ 5.2 Model Architectures
Once the model is trained on large potential data, we extract the intermediate knowledge $\mathbf{a}_{s}^{n}$, $\mathbf{a}_{v}^{n}$, $\mathbf{a}_{t}^{n}$, $\mathbf{b}_{s}^{n}$, and $\mathbf{b}_{v}^{n}$, from a pretrained PFP-$\mathrm{L}_{n}$ as the input for our equivariant readout layers.
@@
Overview The input materials are preprocessed to identify appropriate neighbor atoms under the periodic boundary conditions for graph construction. Each material is represented by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The initialized atom attributes contain three types of information with different ranks: $\mathcal{V}=\{(a_{s}^{(0)},a_{v}^{(0)},a_{t}^{(0)})_{i}\}_{i=1:N^{a}}$, where $a_{s}^{(0)}\in\mathbb{R}^{C_{s}^{a}}$, $a_{v}^{(0)}\in\mathbb{R}^{C_{v}^{a}\times 3}$, $a_{t}^{(0)}\in\mathbb{R}^{C_{t}^{a}\times 3\times 3}$, and $N^{a}$ is the number of atoms. $a_{s}$ is initialized by looking up the table, in which each element is mapped to a high-dimension encoding with summation equal to the atomic number divided by 2. Both atom vector $a_{v}$ and atom tensor $a_{t}$ are initialized to be zeros. Two types of bond attributes are prepared: =$\mathcal{E}=\{(b_{s}^{(0)},b_{v}^{(0)})_{k}\}_{k=1:N^{b}}$, in which $b_{s}^{(0)}\in\mathbb{R}^{C_{s}^{b}}$, $b_{v}^{(0)}\in\mathbb{R}^{C_{v}^{b}\times 3}$, $N^{b}$ is the number of bonds between atom pairs within the cutoff radius $R_{c}$, and $C_{s}^{a}$, $C_{v}^{a}$, $C_{t}^{a}$, $C_{s}^{b}$, $C_{v}^{b}$ are the corresponding number of channels for each feature. Both $b_{v}^{(0)}$ and $b_{s}^{(0)}$ are filled with zeros initially in this work.
@@
TeaNet[31] is used as the base GNN architecture in PFP[32, 33]. The TeaNet architecture incorporates a second-order Euclidean tensor for high-order geometry interaction and performs equivariant graph convolutions using its information. The original material[31] provides a step-by-step explanation for TeaNet implementation, please check it for more details. PFP has several modifications on TeaNet, such as introducing the Morse-style two-body potential term and so on, see the corresponding material[32] for details. PFP hold the scalar features $a_{s}$, $b_{s}$ invariant and the vector/tensor features $a_{v}$, $a_{t}$, $b_{v}$ equivariant to 3D rotations. PFP calculates local interactions in each of its GCN layers. 5-layer GCN layers are constructed with the cutoff radius $R_{c}$ as 3, 3, 4, 6 and 6Å, respectively. In other words, latent embeddings in the earlier convolutional layers hold less distant information. As a result, the maximum distance of atomic interactions counted in PFP is 22Å after the 5-layer information propagation. The intermediate atom/bond features $a_{s}^{(l)}$, $a_{v}^{(l)}$, $a_{t}^{(l)}$, $b_{s}^{(l)}$, and $b_{v}^{(l)}$ extracted from the $l$-th GCN layer are used as the input to the designed DTNN which contains $N$ equivariant blocks and a readout neural network.
@@
$\mathcal{L}^{k}:x\mapsto\mathbf{W_{k}}x$ (8)
@@
$\mathcal{L}_{g}^{k}:x\mapsto g(\mathbf{W_{k}}x+\mathbf{b_{k}})$ (9)
where $\mathbf{W_{k}}$ and $\mathbf{b_{k}}$ are learnable parameters, $g(x)$ is the activation function which can be substituted by $s(x)$ as the SiLU activation function and $\sigma(x)$ as the sigmoid activation function. The SiLU activation function $s(x)$[55, 56, 57] is defined by the sigmoid function $\sigma(x)$ multiplied by its input $x$
@@
Equivariant block The equivariant block takes $a_{s}^{(l,n)}$, $a_{v}^{(l,n)}$, $a_{t}^{(l,n)}$, $b_{s}^{(l)}$, and $b_{v}^{(l)}$ as inputs, where $l$ denotes the GCN layer number of PFP, and $n$ denotes the layer number of the equivariant block. The perceptron models are applied to the atom scalar features $a_{s}^{(l,n)}$ and edge scalar features $b_{s}^{(l)}$, respectively,
$a_{s1}^{(l,n)}=\xi_{a_{s}1}^{K}(a_{s}^{(l,n)})$ (13)
$b_{s1}^{(l,n)}=\xi_{b_{s}1}^{K}(b_{s}^{(l,n)})$ (14)
To keep the equivariance of vector features and tensor features, the linear transformations without the bias term are applied to the atom vector features $a_{v}^{(l,n)}$, atom tensor features $a_{t}^{(l,n)}$, and bond vector features $b_{v}^{(l)}$.
$a_{v1}^{(l,n)}=\mathcal{L}_{a_{v}1}(a_{v}^{(l,n)})$ (15)
$a_{t1}^{(l,n)}=\mathcal{L}_{a_{t}1}(a_{v}^{(l,n)})$ (16)
$b_{v1}^{(l)}=\mathcal{L}_{b_{v}1}(b_{v}^{(l)})$ (17)
The neighbor set of atom $i$ is denoted as $\mathcal{N}_{i}$. Let $a_{s1,i}^{(l,n)}$ denote the scalar feature of atom $i$, and $b_{s1,\{i,j\}}^{(l)}$ denote the scalar feature of the bond connecting atom $i$ and atom $j$. Other vector and tensor features are represented in a similar manner. The bond scalar features $b_{s1}^{(l)}$ and vector features $b_{v1}^{(l)}$ are aggregated to atom scalar features $a_{s1}^{(l,n)}$ and vector features $a_{v1}^{(l,n)}$ accordingly, thus for each atom $i$
$a_{s2,i}^{(l,n)}=a_{s1,i}^{(l,n)}+\sum_{j\in\mathcal{N}_{i}}b_{s1,\{i,j\}}^{(l)}$ (18)
$a_{v2,i}^{(l,n)}=a_{v1,i}^{(l,n)}+\sum_{j\in\mathcal{N}_{i}}b_{v1,\{i,j\}}^{(l)}$ (19)
Then the norm of node vector features and tensor features are combined with the node scalar features and fed into another perceptron model.
$a_{s3}^{(l,n)}=\xi_{a_{s}3}^{K}(\mathrm{concat}(a_{s2}^{(l,n)}+\mathrm{norm}(a_{v2}^{(l,n)})+\mathrm{norm}(a_{t2}^{(l,n)}))$ (20)
@@
$a_{s4}^{(l,n+1)},a_{v3}^{(l,n)},a_{v3}^{(l,n)}=\mathrm{split}(a_{s3}^{(l,n)})$ (21)
$a_{v4}^{(l,n+1)}=\sigma(a_{v3}^{(l,n)})\odot\mathcal{L}_{a_{v}2}(a_{v}^{(l,n)})$ (22)
$a_{t4}^{(l,n+1)}=\sigma(a_{t3}^{(l,n)})\odot\mathcal{L}_{a_{t}2}(a_{t}^{(l,n)})$ (23)
Finally, the updated $a_{s4}^{(l,n+1)}$, $a_{v4}^{(l,n+1)}$ and $a_{t4}^{(l,n+1)}$ are used as the input for the next equivariant readout block.
Readout The final layer output, i.e., layer $N$, of the equivariant blocks $a_{s}^{(l,N)}$, $a_{v}^{(l,N)}$ and $a_{t}^{(l,N)}$ are aggregated for the tensorial property prediction. Since the dielectric constant is an intensive property, we take the average aggregation of the atom features in the graph representation of materials. The gated MLP applied to the scalar features $a_{s}^{(l,N)}$ is found to help improving the accuracy compared with the normal MLP. Therefore, the atom scalar features is processed by
$a_{s}^{\prime}=\phi^{K}(a_{s}^{(l,N)})$ (24)
@@
$a_{v}^{\prime}=\mathcal{L}_{a_{v}3}(a_{v}^{(l,N)})$ (25)
$a_{t}^{\prime}=\mathcal{L}_{a_{t}3}(a_{t}^{(l,N)})$ (26)
@@
$\boldsymbol{\epsilon}=\frac{1}{N}\sum^{N}(a_{s}^{\prime}\mathbf{I}_{3}+a_{v}^{\prime}\otimes a_{v}^{\prime}+a_{t}^{\prime})$ (27)
where $\mathbf{I}_{3}$ is the $3\times 3$ tensor with ones on the diagonal and zeros elsewhere, and $\otimes$ represents the outer product operation. The first models isotropic, and the final two terms add the anisotropic components of the dielectric tensor. Therefore, the final output of our model is the $3\times 3$ dielectric tensor following the equivariance of the materials input.

### v2

@@ 2.4.2 High-dielectric materials
Figure 8: (a-b) The joint distribution of band gaps $E_{g}$ and electronic dielectric constants $\varepsilon^{\infty}$ (a) or ionic dielectric constants $\varepsilon^{0}$ (b) for 6,648 training structures and 14,375 stable candidate structures. $\varepsilon^{\infty}$ and $\varepsilon^{0}$ of training structures and $E_{g}$ of all structures are obtained from computational results in MP, while $\varepsilon^{\infty}$ and $\varepsilon^{0}$ of candidate structures are predicted by DTNet-$\varepsilon^{\infty}/\varepsilon^{0}$. The 5-level densities of data points are drawn in contours using the kernel density estimation for structures both in the candidate set and original set. (c) DFT-calculated $E_{g}-\varepsilon$ data points colored by their figure of merit defined by $E_{g}\cdot\varepsilon$ for 4,146 stable structures in training data and 35 new stable structures screened by DTNet across 3 iterations. (d) Structure visualization of the top 3 materials by FOM as validated by DFPT calculations.
2.4.3 Highly anisotropic dielectrics
Crystalline dielectrics can exhibit anisotropic dielectric properties w.r.t. the direction of the applied electric field due to their cell structures and atomic arrangements. Crystals with highly anisotropic dielectric properties can demonstrate unique characteristics, such as enhanced piezoelectricity, ferroelectricity [47], and optical performance [48]. Lou et al. explored the discovery of crystals with high anisotropy in high-frequency dielectric tensors, focusing solely on the electronic contribution [35]. In this task, we tackle the more challenging task of discovering highly anisotropic static dielectric tensors, incorporating both electronic and ionic contributions, to achieve significantly greater anisotropy under the static external electric field.
Given the limited number of training samples with high anisotropy, we conducted a 3-round active learning for virtual screening using the same strategy to more effectively identify marginal candidates exhibiting high anisotropic ratio defined by $\alpha_{r}=\lambda_{\textrm{max}}/\lambda_{\textrm{min}}$ [35]. All molecular crystalline candidates were excluded prior to virtual screening due to their typically high anisotropy and potential instability under the conditions of interest. For example, we evaluated the molecular crystal BrCl (mp-1066781). Despite exhibiting a high total anisotropy of $\alpha_{r}=26.774$, its melting point is as low as $-66^{\circ}C$, making it unstable under room temperature conditions. Consequently, 41 out of the 60 screened structures were successfully validated via DFPT computations. Fig. 9(a) presents the distribution of anisotropic ratios $\alpha_{r}$. Notably, only 17 structures in the training dataset exhibit $\alpha_{r}$ exceeding 10, with the highest of 35.027. In contrast to the average $\alpha_{r}$ of 1.665 observed in stable crystals within the dataset, the three iterative rounds produced significantly improved average $\alpha_{r}$ of 5.608, 12.946, and 17.687, respectively. The top three DFPT-validated candidates were identified during the second and third iterations, with their structures visualized in Fig. 9(b). Notably, CsZrCuSe3 and BaNiO3 are 3D materials, which may offer greater robustness and suitability for industrial applications compared to the 1D material SeI2. The ionic components significantly contribute to the anisotropy in all three materials. For CsZrCuSe3, the eigenvalue ranges for the electronic, ionic, and total contributions are [5.152, 12.851], [4.306, 1141.694], and [9.458, 1152.884], respectively, achieving electronic, ionic and total ratios of 2.494, 265.140 and 128.890, respectively. The anisotropic ratios for SeI2 are 2.745 (electronic), 316.090 (ionic), and 96.763 (total), and for BaNiO3 they are 1.749 (electronic), 92.904 (ionic), and 61.026 (total). Compared to the largest $\alpha_{r}=35.027$ in the training set (BrNO3, mp-29526), CsZrCuSe3 is 93.863 larger and 2.680 times greater. A recent research has also highlighted the excellent optoelectronic and thermoelectric properties of CsZrCuSe3 [49], underscoring its great potential for practical applications.
Figure 9: (a) Comparison of anisotropic distributions of 4,146 stable structures in training data and 41 new stable structures screened by DTNet over 3 iterations. (d) Structure visualization of the top 3 materials by anisotropic ratios as validated by DFPT calculations.
@@ 3 Discussion
Transfer learning is the most popular approach to tackle the data bottleneck which is common in materials science. Previous studies have utilized transfer learning in GNN models, both for the same properties and across different properties, to demonstrate enhanced accuracy in materials science [50, 51, 15]. For example, Chen et al. developed the AtomSets framework [52], which employs a straightforward model architecture to process structure-embedded scalar information read from a pretrained MEGNet model [26], enabling high accuracy predictions with limited datasets. However, the feasibility of transfer learning for properties across different tensorial orders has been less explored. Our study indicates that more than scalar features, it is also possible to transfer higher-rank equivariant compositional and structural information from a pretrained EGNN energy model to a different tensorial property prediction with a small amount of available data, while maintaining the equivariance. These equivariant representations may hold shared structural or even electronic-level information, since the PFP model aims to emulate iterative electronic relaxation. This is evident through the accurate predictions our model achieved for both electronic and ionic dielectric constants, which are produced by different microscopic mechanisms. Furthermore, the pretrained PFP serves as an encoder that integrates electron density information, assisting in comprehending the electronic-level relationship between the $E_{g}$ and $\varepsilon^{\infty}$. Apart from dielectric constants, several works attempted to predict tensorial properties of materials such as the polarizability [53], quadrupole [54], by using equivariant architectures. The accuracy of such tasks could potentially be improved by fixing parameters in the shallower layers of an EGNN model pretrained on commonly computed properties like energy.
@@
The PFP model is pretrained on over 22 million structures covering the diverse chemical space. It is efficient and expressive to take a a pretrained PFP as the parent model to generate locally-interacted features as the input to a much simpler child model. In the current readout architecture, it consists of only 0.6 million trainable parameters for the downstream tasks of dielectric constant prediction. The model performing best on the validation accuracy appears at the average epoch of 263 out of 5 training runs for total dielectric constants. By stopping the training after 200 epochs without improvement in validation accuracy, the model training can be efficiently completed within approximately 2 hours using about 6GB memory on an Nvidia Tesla-V100 GPU. We proved that such efficient training on encodings from the parent model can bring a promising property predictor even on a small data region as demonstrated in Table 2. The incorporation of second-order equivariant features in our model may also contribute to high data efficiency as the symmetry of the dielectric tensors is automatically guaranteed due to the model architecture. Based on only 6.6k sample structures with labeled tensor containing up to 6 independent elements from MP, our model achieved a mean tensor prediction error of $0.368$ for $\boldsymbol{\varepsilon}^{\infty}$, $1.677$ for $\boldsymbol{\varepsilon}^{0}$ and $1.911$ for $\boldsymbol{\varepsilon}$.
@@
Thanks to widely-supported elements in the pretrained PFP, our model can also give relatively good prediction on rare earth elements, e.g., the Sm element in Ba2SmTaO6. This compound emerged as one of the top high-dielectric candidates identified from a dataset of over 14,000 structures, based directly on its structural graph. The rare earth elements have diverse industrial applications such as glass, lights, magnets, batteries, and catalytic converters and so on [55], due to their unique physical and chemical properties. Our model has the potential to make predictions about properties concerning rare elements, which typically suffer from limited data availability.
@@
In this study, we demonstrate the applicability of our model to two design challenges, utilizing different metrics derived from the predicted dielectric tensors with high accuracy. Accurate predictions of dielectric constants as tensors can contribute to understanding material behavior in the presence of electric fields and their interactions with other materials. This knowledge can facilitate the development of new materials with tailored properties for specific applications, such as energy storage, signal transmission, electronic component design, and cutting-edge fields such as dark matter detection [56].
@@ 4 Conclusion
An equivariant model is introduced to predict the second-order tensorial property, dielectric constants, for inorganic materials across 72 supported elements. By leveraging transfer learning, the PFP was treated as the frozen base model to encode material graphs with integrated elemental and structural information. A light-weight trainable equivariant readout module is connected to integrate information in different tensorial orders for the final outputs. It was discovered that embedded features from deeper GCN layers in PFP contribute to better accuracy in predicting total dielectric constants. After efficiently training the equivariant readout module, our DTNet model achieved superior results on the dielectric task in Matbench, and also higher accuracy in predicting the electronic, ionic, total dielectric tensors compared with state-of-the-art models, PaiNN, M3GNet and MatTen, across various crystal system types. To assess the model’s capability in discovering novel dielectric materials, we utilized it to identify high-score candidates near the tail of the training distribution for two distinct design targets by conducting virtual screening on total static dielectric constants. For applications requiring both high dielectric constants and large band gaps in microelectronic manufacturing, the model successfully identified the top three dielectric materials, Cs2Ti(WO4)3 (mp-1226157, $E_{g}=2.83\mathrm{eV},\varepsilon=180.89$), RbNbWO6 (mp-1219587, $E_{g}=2.93\mathrm{eV},\varepsilon=155.57$) and Ba2SmTaO6 (mp-1214622, $E_{g}=3.36\mathrm{eV},\varepsilon=93.81$), from a comprehensive pool of over 14,000 candidates, with 35 DFPT calculations performed. For dielectrics with high anisotropy, the top three materials (including two 3D materials) identified by the model exhibit high anisotropic ratios of 128.890 (CsZrCuSe3, mp-7152), 96.763 (SeI2, mp-861871) and 61.026 (BaNiO3, mp-19138), respectively, validated by 41 DFPT calculations. Promising candidates that significantly outperform the training distribution were successfully identified in both tasks.
@@ 5.2 Model Architectures
Once the model is trained on large potential data, we extract the intermediate knowledge $\boldsymbol{a}_{s}^{n}$, $\boldsymbol{a}_{v}^{n}$, $\boldsymbol{a}_{t}^{n}$, $\boldsymbol{b}_{s}^{n}$, and $\boldsymbol{b}_{v}^{n}$, from a pretrained PFP-$\mathrm{L}_{n}$ as the input for our equivariant readout layers.
@@
Overview The input materials are preprocessed to identify appropriate neighbor atoms under the periodic boundary conditions for graph construction. Each material is represented by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The initialized atom attributes contain three types of information with different ranks: $\mathcal{V}=\{(\boldsymbol{a}_{s}^{0},\boldsymbol{a}_{v}^{0},\boldsymbol{a}_{t}^{0})_{i}\}_{i=1:N^{a}}$, where $\boldsymbol{a}_{s}^{0}\in\mathbb{R}^{C_{s}^{a}}$, $\boldsymbol{a}_{v}^{0}\in\mathbb{R}^{C_{v}^{a}\times 3}$, $\boldsymbol{a}_{t}^{0}\in\mathbb{R}^{C_{t}^{a}\times 3\times 3}$, and $N^{a}$ is the number of atoms. $\boldsymbol{a}_{s}^{0}$ is initialized by looking up a predefined table, in which each element is mapped to a high-dimension encoding with summation equal to the atomic number divided by 2. Both atom vector $\boldsymbol{a}_{v}^{0}$ and atom tensor $\boldsymbol{a}_{t}^{0}$ are initialized to be zeros. Two types of bond attributes are prepared: =$\mathcal{E}=\{(\boldsymbol{b}_{s}^{0},\boldsymbol{b}_{v}^{0})_{k}\}_{k=1:N^{b}}$, in which $\boldsymbol{b}_{s}^{0}\in\mathbb{R}^{C_{s}^{b}}$, $\boldsymbol{b}_{v}^{0}\in\mathbb{R}^{C_{v}^{b}\times 3}$, $N^{b}$ is the number of bonds between atom pairs within the cutoff radius $R_{c}$, and $C_{s}^{a}$, $C_{v}^{a}$, $C_{t}^{a}$, $C_{s}^{b}$, $C_{v}^{b}$ are the corresponding number of channels for each feature. Both $\boldsymbol{b}_{v}^{0}$ and $\boldsymbol{b}_{s}^{0}$ are filled with zeros initially in this work.
@@
TeaNet [32] is used as the base GNN architecture in PFP [33, 34]. The TeaNet architecture incorporates a second-order Euclidean tensor for high-order geometry interaction and performs equivariant graph convolutions using its information. The original material [32] provides a step-by-step explanation for TeaNet implementation for more details. PFP has several modifications on TeaNet, such as introducing the Morse-style two-body potential term and so on, see the corresponding material [33] for details. PFP holds the scalar features $\boldsymbol{a}_{s}$, $\boldsymbol{b}_{s}$ invariant and the vector/tensor features $\boldsymbol{a}_{v}$, $\boldsymbol{a}_{t}$, $\boldsymbol{b}_{v}$ equivariant to rotations/inversions. PFP calculates local interactions in each of its GCN layers. 5-layer GCN layers are constructed with the cutoff radius $R_{c}$ as 3, 3, 4, 6 and 6Å, respectively. In other words, latent embeddings in the earlier convolutional layers hold less distant information. As a result, the maximum distance of atomic interactions counted in PFP is 22Å after the 5-layer information propagation. The intermediate atom/bond features $\boldsymbol{a}_{s}^{n}$, $\boldsymbol{a}_{v}^{n}$, $\boldsymbol{a}_{t}^{n}$, $\boldsymbol{b}_{s}^{n}$, and $\boldsymbol{b}_{v}^{n}$ extracted from the $n$-th GCN layer are used as the input to the designed DTNet which contains $M$ equivariant blocks and a readout neural network.
@@
$\mathcal{L}^{k}:x\mapsto\boldsymbol{W_{k}}x$ (8)
@@
$\mathcal{L}_{g}^{k}:x\mapsto g(\boldsymbol{W_{k}}x+\boldsymbol{b_{k}})$ (9)
where $\boldsymbol{W_{k}}$ and $\boldsymbol{b_{k}}$ are learnable parameters, $g(x)$ is the activation function which can be substituted by $s(x)$ as the SiLU activation function and $\sigma(x)$ as the sigmoid activation function. The SiLU activation function $s(x)$ [62, 63, 64] is defined by the sigmoid function $\sigma(x)$ multiplied by its input $x$
@@
Equivariant block The equivariant block takes $\boldsymbol{a}_{s}^{(n,m)}$, $\boldsymbol{a}_{v}^{(n,m)}$, $\boldsymbol{a}_{t}^{(n,m)}$, $\boldsymbol{b}_{s}^{n}$, and $\boldsymbol{b}_{v}^{n}$ as inputs, where $n$ denotes the GCN layer number in PFP, and $m$ denotes the layer number of the equivariant block. The perceptron models are applied to the atom scalar features $\boldsymbol{a}_{s}^{(n,m)}$ and edge scalar features $\boldsymbol{b}_{s}^{n}$, respectively,
$\boldsymbol{a}_{s1}^{(n,m)}=\xi_{a_{s}1}^{K}(\boldsymbol{a}_{s}^{(n,m)})$ (13)
$\boldsymbol{b}_{s1}^{(n,m)}=\xi_{b_{s}1}^{K}(\boldsymbol{b}_{s}^{(n,m)})$ (14)
To keep the equivariance of vector features and tensor features, the linear transformations without the bias term are applied to the atom vector features $\boldsymbol{a}_{v}^{(n,m)}$, atom tensor features $\boldsymbol{a}_{t}^{(n,m)}$, and bond vector features $\boldsymbol{b}_{v}^{n}$.
$\boldsymbol{a}_{v1}^{(n,m)}=\mathcal{L}_{a_{v}1}(\boldsymbol{a}_{v}^{(n,m)})$ (15)
$\boldsymbol{a}_{t1}^{(n,m)}=\mathcal{L}_{a_{t}1}(\boldsymbol{a}_{t}^{(n,m)})$ (16)
$\boldsymbol{b}_{v1}^{n}=\mathcal{L}_{b_{v}1}(\boldsymbol{b}_{v}^{n})$ (17)
The neighbor set of atom $i$ is denoted as $\mathcal{N}_{i}$. Let $\boldsymbol{a}_{s1,i}^{(n,m)}$ denote the scalar feature of atom $i$, and $\boldsymbol{b}_{s1,\{i,j\}}^{n}$ denote the scalar feature of the bond connecting atom $i$ and atom $j$. Other vector and tensor features are represented in a similar manner. The bond scalar features $\boldsymbol{b}_{s1,\{i,j\}}^{n}$ and vector features $\boldsymbol{b}_{v1,\{i,j\}}^{n}$ are aggregated to atom scalar features $\boldsymbol{a}_{s1,i}^{(n,m)}$ and vector features $\boldsymbol{a}_{v1,i}^{(n,m)}$ accordingly, thus for each atom $i$
$\boldsymbol{a}_{s2,i}^{(n,m)}=\boldsymbol{a}_{s1,i}^{(n,m)}+\sum_{j\in\mathcal{N}_{i}}\boldsymbol{b}_{s1,\{i,j\}}^{n}$ (18)
$\boldsymbol{a}_{v2,i}^{(n,m)}=\boldsymbol{a}_{v1,i}^{(n,m)}+\sum_{j\in\mathcal{N}_{i}}\boldsymbol{b}_{v1,\{i,j\}}^{n}$ (19)
[…]
$\boldsymbol{a}_{s3}^{(n,m)}=\xi_{a_{s}3}^{K}(\mathrm{concat}(\boldsymbol{a}_{s2}^{(n,m)}+\|\boldsymbol{a}_{v2}^{(n,m)}\|+\|\boldsymbol{a}_{t2}^{(n,m)}\|)$ (20)
@@
$\boldsymbol{a}_{s4}^{(n,m+1)},\boldsymbol{a}_{v3}^{(n,m)},\boldsymbol{a}_{v3}^{(n,m)}=\mathrm{split}(\boldsymbol{a}_{s3}^{(n,m)})$ (21)
$\boldsymbol{a}_{v4}^{(n,m+1)}=\sigma(\boldsymbol{a}_{v3}^{(n,m)})\odot\mathcal{L}_{a_{v}2}(\boldsymbol{a}_{v}^{(n,m)})$ (22)
$\boldsymbol{a}_{t4}^{(n,m+1)}=\sigma(\boldsymbol{a}_{t3}^{(n,m)})\odot\mathcal{L}_{a_{t}2}(\boldsymbol{a}_{t}^{(n,m)})$ (23)
Finally, the updated $\boldsymbol{a}_{s4}^{(n,m+1)}$, $\boldsymbol{a}_{v4}^{(n,m+1)}$ and $\boldsymbol{a}_{t4}^{(n,m+1)}$ are used as the input for the next equivariant readout block.
Readout The final layer output, i.e., layer $M$, of the equivariant blocks $\boldsymbol{a}_{s}^{(n,M)}$, $\boldsymbol{a}_{v}^{(n,M)}$ and $\boldsymbol{a}_{t}^{(n,M)}$ are aggregated for the tensorial property prediction. Since the dielectric constant is an intensive property, we take the average aggregation of the atom features in the graph representation of materials. The gated MLP applied to the scalar features $\boldsymbol{a}_{s}^{(n,M)}$ is found to help improving the accuracy compared with the normal MLP. Therefore, the atom scalar features is processed by
$\boldsymbol{a}_{s}^{\prime}=\phi^{K}(\boldsymbol{a}_{s}^{(n,M)})$ (24)
@@
$\boldsymbol{a}_{v}^{\prime}=\mathcal{L}_{a_{v}3}(\boldsymbol{a}_{v}^{(n,M)})$ (25)
$\boldsymbol{a}_{t}^{\prime}=\mathcal{L}_{a_{t}3}(\boldsymbol{a}_{t}^{(n,M)})$ (26)
@@
$\boldsymbol{\varepsilon}=\frac{1}{N^{a}}\sum^{N^{a}}(\boldsymbol{a}_{s}^{\prime}\boldsymbol{I}_{3}+\boldsymbol{a}_{v}^{\prime}\otimes\boldsymbol{a}_{v}^{\prime}+\boldsymbol{a}_{t}^{\prime})$ (27)
where $\boldsymbol{I}_{3}$ is the $3\times 3$ identity tensor, and $\otimes$ represents the outer product operation. The first models isotropic, and the final two terms add the anisotropic components of the dielectric tensor. Therefore, the final output of our model is the $3\times 3$ dielectric tensor following the equivariance of the materials input.